OpenStudy (anonymous):

Solve the equation 8= √2 ^(4x+8) . Using complete sentences, explain the strategy used in solving this equation.

OpenStudy (harsimran_hs4):

take log of both the sides log 8 = log (sqrt(2)^(4x + 8)) ...can you proceed now ?

OpenStudy (nurali):

\[8=\sqrt{2^{4x+8}}\] Squaring Both Sides \[(8)^2=(\sqrt{2^{(4x+8)}})^2\] \[64=2^{(4x+8)}\] \[2^6=2^{(4x+8)}\] \[6=4x+8\] \[4x=6-8\] \[4x=-2\] \[x=-\frac{ 1 }{ 2 }\]